OpenStudy (anonymous):

Smith found out that 15 out of 105 students in his chemical course failed the first exam. What fractional part of the class failed the exam. What fractional part of the class period?

OpenStudy (anonymous):

Fraction of the class that failed = 15/105 = 3/21 = 1/7 Fraction of the class that passed = 90/105 = 18/21 = 6/7